# thanks: Kavli IPMU Fellow. Corresponding author

### v1

@@ Appendix A Density power spectrum
Where the extra $\sqrt{2}$ in the argument of the exponential comes from the convolution in $\rho\propto|\psi|^{2}$. We have used the fact that $v=\hbar k$. Note that the power spectrum in our simulations is consistent with previous similar work Dalal et al. (2021), see Figure 10.
Appendix B Spatial correlations
Figure 11: Here we show the qualitative behavior of spatial correlations of observables in our ultralight dark matter simulations. The left panel shows a density slice. We can clearly see that there exist granules and spatial correlation in the density on the scale of the de Broglie wavelength which here is $\lambda_{\mathrm{db}}\sim 0.1\,\mathrm{kpc}$. The middle plot shows the Shapiro delay from the center of the simulation to each point evaluated everywhere in this same slice. We see that correlations are projected out from the center, which here is the position of the observer. In the right panel, we plot the redshift at each point in the slice. We see that correlations are larger than in the density, this is because Poisson’s equation amplifies large scale modes in the potential. The scale on the color-bars in each subplot is arbitrary. The field mass here is $m=10^{-22}\,\mathrm{eV}$.
@@ Appendix B Spatial correlations
In this appendix, we discuss the spatial correlations in the system qualitatively. We first do this by taking a snapshot of a simulation and plotting the density, Shapiro delay, and redshift everywhere in a slice, where the observer is in the center of the slice. This is plotted in Figure 11. In the left panel, we can see the characteristic granule and spatial correlations on the scale of the de Broglie wavelength, which for the $m=10^{-22}\,\mathrm{eV}$ field simulated in Figure 11, is $\lambda_{\mathrm{db}}\sim 0.1\,\mathrm{kpc}$. In the right panel, we see the spatial gravitational potential. We can see that it exhibits correlations on larger scales than the density. This corroborates our expectations as the Poisson equation amplified large-scale modes, see equation (19). In the middle panel, we have plotted the Shapiro delay at each position in the slice. We can see that the correlations here are more angular instead of spatial, i.e. correlations go along projections out from the center. This is because the Shapiro delay involves an integral along the line of sight, meaning that points will be correlated with other points along the same line of sight.
We plot the time averaged spatial and angular correlations in Figure 12. We can see that the correlation has the expected decay on the de Brolie scale.
Appendix C Other ULDM effects on pulsar timing
@@
C.1 Compton scale redshift
@@
$\Phi(x,t)=\Phi_{0}(x,t)+\pi\frac{G\rho}{m^{2}}\cos{(\omega_{c}t+2\alpha(x))}\,,$ (27)
@@
C.2 Doppler shift
The oscillating de Broglie scale potential provides accelerations to stars in orbit in an ultralight dark matter halo. These accumulated accelerations create a random peculiar velocity between the Earth and any given pulsar. This velocity creates a stochastic Doppler shift in the observed pulsar frequency which varies on the de Broglie timescales.
The spectrum and amplitude of this signal were first worked out in Kim and Mitridate (2024). They used the quasi-particle approximation to argue that the average acceleration felt by a star in orbit due to a granule should be approximately
$\bar{a}\sim\frac{Gm_{\mathrm{eff}}}{\lambda^{2}}\,.$ (29)
This creates a Doppler shift over a de Broglie time of $z^{D}\sim\delta v/c\sim\bar{a}\tau/c$. These kicks accumulate randomly and result in a total number of kicks $N_{\mathrm{enc}}/\tau\sim T_{\mathrm{obs}}/\tau$ where $T_{\mathrm{obs}}$ is the total observation time. The integral of this frequency change then gives an expected time delay signal described by equation (1) in Kim and Mitridate (2024)
$\delta t^{D}$ $\sim(\bar{a}\tau^{2})\left(\frac{T_{\mathrm{obs}}}{\tau}\right)^{3/2}\,,$
$\sim 0.1\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{3/2}\left(\frac{\rho}{0.4\,\mathrm{GeV/cm^{3}}}\right)\,\mathrm{ns}\,.$ (30)

### v2

@@ Appendix B Density power spectrum
Where the extra $\sqrt{2}$ in the argument of the exponential comes from the convolution in $\rho\propto|\psi|^{2}$. We have used the fact that $v=\hbar k/m$. Note that the power spectrum in our simulations is consistent with previous similar work Dalal et al. (2021), see Figure 11.
Appendix C Spatial correlations
@@
In this appendix, we discuss the spatial correlations in the system qualitatively. We plot the time averaged spatial and angular correlations of a simulation in Figure 12. We can see that the correlation has the expected decay on the de Broglie scale.
Appendix D Other ULDM effects on pulsar timing
@@
D.1 Compton scale redshift
@@
$\Phi(\vec{x},t)=\Phi_{0}(\vec{x},t)+\pi\frac{G\rho}{m^{2}}\cos{(\omega_{c}t+2\alpha(\vec{x}))}\,$, (47)